0:35 Powering up the WFS. Noisy OCAM. Cycling of BCU2k resolved.

5:14 Collimating and pointing for 1st (and only ESM target), GCS reporting 0.4-0.5"

SpARCS Tail. AOref catalog magnitude 13.53. Some clouds tonight. Seeing on the DIMM 0.9"

5:18 Preset to SpARCS tail. Seeing 0.45=0.5" on the guider.

05:22 OCAM not live during Center pupils caused adsec failure. Executing adsc_stop/start. Restarted the adsec arbitrator for good measure.

Flat set with 635 modes

Josh was having some issues releasing presets and with PMC:incluence functions on the left primary.

5:45 Resending preset to SpARCS tail. Seeing 0.55" on the guiders.

Camera lens is saturated with a rotator angle of 34.

The rotator is set to a specific rotator position for the science and is not adjustable. We are going to try and come at this position from a different rotator angle. We also note that the rerotator tracking is out of range. Josh is setting the rotator angle to 400 and we are resending the preset.

5:57 Resending preset. Camera lens is still saturated in Y

This program had a specific PA with a blind offset and a 180 offset would not have worked

No more.

-- JenniferPower - 27 Mar 2022
Topic revision: r1 - 27 Mar 2022, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback